# Using Large Language Models for Automatic Hint Ranking and Generation

### v1

@@ 5.4 HintRank Evaluation Method
Which hint is better to find the answer of this question: $q$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between "Hint_1" and "Hint_2" without any explanations.
@@
Which hint is better to find the answer of this question: $q$. The answer for this question is $a$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between "Hint_1" and "Hint_2" without any explanations.
@@
We benchmark HintRank against the Convergence metric which turned out to be useful for hint ranking assessment as indicated in Figure 3. To convert pairwise rankings to listwise rankings, we apply the Bradley–Terry model Bradley and Terry (1952). We evaluate the correlation between the rankings with Pearson Correlation Mining (2006).
@@
Table 5 outlines the key features and differences among various scenarios. The results indicate that with the increase in the size and power of LLMs, both accuracy and correlation improve. Additionally, the answer-aware approach yields better outcomes compared to the answer-agnostic method, suggesting that the presence of an answer enables LLMs to evaluate hints more effectively. Moreover, finetuned versions outperform their vanilla counterparts, demonstrating that the WikiHint dataset is well-suited for model fine-tuning to rank hints. Surprisingly, the BERT-base method outperforms LLMs, including finetuned versions in the answer-aware scenario. This holds true for both Bert-FTwoA and Bert-FTwA, although BERT-base also performs better in the answer-aware approach compared to answer-agnostic. BERT-base methods achieve higher accuracy than LLMs and convergence, but in terms of correlation, LLaMA-3.1-70b exhibits the best performance. The effectiveness of BERT-base methods may be attributed to the strengths of encoder-based models like BERT in classification tasks over decoder-based models. Utilizing BERT-based models instead of LLMs enhances the speed and accessibility of HintRank, reducing computational demands. Figure 13 in Appendix A shows that accuracy improves as the rank difference between hints increases, indicating it’s harder to correctly order hints with closer ranks.
6 Conclusions
@@
In this paper, we introduced the first manually created dataset for hint generation and hint ranking. We also presented a new lightweight method for evaluating and ranking hints. To demonstrate the effectiveness of our dataset, we conducted experiments where humans attempted to answer questions with and without the use of hints. The results confirm that the hints are of sufficient quality to assist users. We then finetuned LLMs using our dataset, prompting them to generate new hints for different questions. The high quality of the generated hints indicates that our dataset is well-suited for finetuning LLMs for HG task. We also finetuned BERT and LLMs on the dataset for the task of hint ranking and evaluated their performance. The results reveal that encoder-based models outperform decoder-based models in hint ranking.
@@
In future, we plan to generate personalized hints tailored to the knowledge of askers. The main challenge here will be to develop appropriate datasets and solutions for user profiling.
Limitations
Our study has the following limitations:
  • The need for generative capabilities in hint generation task necessitates the use of LLMs. However, this dependency is a limitation as fine-tuning and prompting LLMs require extensive computational resources and are time-consuming.
  • Our research focus on factoid questions may limit its applicability to other types of questions that involve more complex or abstract answers. Factoid questions, by their nature, provide clear and concrete answers, which simplifies automated hint generation and evaluation but may not fully capture the breadth of human inquiry.
  • The WikiHint dataset is exclusively written in the English language. While this facilitates accessibility for a global audience and ensures compatibility with most existing Large Language Models, it also limits the dataset’s applicability in multilingual or non-English contexts, potentially excluding non-English speakers and diverse linguistic data.
Ethical Considerations
Our study utilizes GPT models, which are covered by the OpenAI License and Apache-2.0 license, and the LLaMA model, which is distributed under Meta’s LLaMA 2 Community License Agreement. We comply with these licensing agreements in all applications. Additionally, the datasets we use are sourced from repositories that are approved for academic use. The artifacts developed during our research are made available under the MIT license to facilitate straightforward modifications and use by the research community. We ensure that our data management, model training, and dissemination practices meet ethical standards and legal requirements associated with each artifact we use.
Appendix A Appendix
Figure 7: The distribution of Train and Test subsets.
Figure 8: Question difficulty based on different question types.
Source ChatGPT NQ SQuAD 2.0
Difficulty 0.43 0.34 0.38
Table 6: Question distributions based on the their sources and difficulty.
Figure 9: The MTurk Worker interface for the hint generation task.
Figure 10: The summarized instructions for the hint generation task.
Figure 11: The detailed instructions for the hint generation task.
Figure 12: Good and bad examples for the hint generation task.
Attribute Description
question The content of the question.
major The major category of the question.
minor The specific sub-category of the question.
entity Content of an entity
ent_type Type of entity (e.g., GPE, PERSON).
start_index Start index of the entity in the question.
end_index End index of the entity in the question.
wikipedia_page_title The title of the corresponding Wikipedia page to the entity in question.
readability Indicates the readability score
familiarity Indicates the familiarity score
difficulty Indicates the difficulty level of the question
Table 7: A detailed description of attributes of a question in WikiHint.
Attribute Description
answer The actual answer.
entity Content of an entity as identified within the answer.
ent_type Type of entity.
start_index Start index of the entity in the answer.
end_index End index of the entity in the answer.
wikipedia_page_title The title of the corresponding Wikipedia page to the entity in the answer.
familiarity Indicates the familiarity score
difficulty Indicates the difficulty level of the answer
Table 8: A detailed description of attributes of an answer in WikiHint.
Attribute Description
hint Hint provided to assist with the question.
source URL source of the hint.
entity Content of entities mentioned in the hint.
wikipedia_page_title The title of the corresponding Wikipedia page to the entities in the hint.
relevance Indicates the relevance score
readability Indicates the readability score
convergence Indicates the convergence score
familiarity Indicates the familiarity score
answer_leakage Indicates the answer leakage score
rank Priority or helpfulness rating of the hint
Table 9: A detailed description of attributes of a hint in WikiHint.
Question What is the driest country in sub-Saharan Africa?
Fifth Best Hint The earliest settlers in this country were in the 18th century and crossed the Orange River to move into the area.
Fourth Best Hint The name of this country is derived from the oldest desert on Earth.
Third Best Hint This country’s coat of arms is a shield with the same design as the flag, 2 antelopes and a red blue and white bird.
Second Best Hint This country is a country in southern Africa with its western border along the Atlantic Ocean.
Best Hint This country’s flag is blue and green with a red stripe down the middle and the symbol of a sun in the top left corner.
Table 10: First example of the WikiHint.
Question What artist received the Presidential Medal of Freedom posthumously in 2018, in recognition of his contributions to American culture?
Fifth Best Hint The artist had karate as a lifelong interest, including its moves in his performances.
Fourth Best Hint The artist served in the Army after he became famous.
Third Best Hint The artist never performed outside North America.
Second Best Hint The artist tried to establish a career in films with pictures such as Jailhouse Rock and Fun in Acapulco.
Best Hint The artist is known for hits such as "Love Me Tender" and "Jailhouse Rock".
Table 11: Second example of the WikiHint.
Question What British comedy team is famous for its ’Four Yorkshiremen’ sketch, portraying exaggerated tales of hardship with humor?
Fifth Best Hint A giant cupid’s foot is repeatedly used in the show.
Fourth Best Hint This British comedy troupe is formed in 1969 consisting of 6 members.
Third Best Hint The team was awarded the AFI Star Award by the American Film Institute in 1998.
Second Best Hint The team received the BAFTA Award for Outstanding British Contribution to Cinema at the 41st British Academy Film Awards in 1988.
Best Hint The Holy Grail and Life of Brian are some of their greatest comedy films.
Table 12: Third example of the WikiHint.
Question What fictional pirate is known for carrying a compass that doesn’t point north but rather to what the user wants most?
Fifth Best Hint One of the pirate’s most famous quotes is "Now, bring me that horizon".
Fourth Best Hint The fictional pirate appeared in the video game series "Kingdom Hearts".
Third Best Hint The fictional pirate has a blood debt to another character called Davy Jones.
Second Best Hint The ship of the fictional pirate is called the Black Pearl.
Best Hint The fictional pirate is played by Johnny Depp in the film series.
Table 13: Fourth example of the WikiHint.
A hint must not include the exact answer explicitly. A hint must be a sentence. A hint must be specific, not generic. A hint must be from the corresponding Wikipedia page. A hint must have a unique rank.
Table 14: A detailed criteria or standards used for verifying the generated hints during the selection process.
Figure 13: Accuracy of HintRank for different hint pairs in different scenarios. The element at position $(r,c)$ represents the accuracy when comparing $\text{Hint}_{1}$ at rank $r$ to $\text{Hint}_{2}$ at rank $c$.

### v2

@@ 5.4. Evaluation of HintRank Method
Which hint is better to find the answer of this question: $q$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between ”Hint_1” and ”Hint_2” without any explanations.
@@
Which hint is better to find the answer of this question: $q$. The answer to this question is $a$. Hint_1: $h_{1}$. Hint_2: $h_{2}$. Just choose between ”Hint_1” and ”Hint_2” without any explanations.
@@
We benchmark HintRank against the Convergence metric which turned out to be useful for evaluating hint ranking as indicated in Figure 8. To convert pairwise rankings to listwise rankings, we apply the Bradley–Terry model (Bradley and Terry, 1952). We then evaluate the correlation between the rankings with the Pearson Correlation (Mining, 2006).
@@
Table 10 outlines the key features and differences between various scenarios. The results indicate that with the increase in the size and power of LLMs, both accuracy and correlation improve. Additionally, the answer-aware approach yields better outcomes compared to the answer-agnostic method, suggesting that the presence of an answer enables LLMs to evaluate hints more effectively. Moreover, finetuned versions outperform their vanilla counterparts, demonstrating that the WikiHint dataset is well-suited for model fine-tuning to rank hints. Surprisingly, the BERT-base method outperforms LLMs, including their finetuned versions in the answer-aware scenario. This holds true for both Bert-FTwoA and Bert-FTwA, although BERT-base also performs better in the answer-aware approach compared to answer-agnostic. BERT-base methods achieve higher accuracy than LLMs and convergence, but in terms of correlation, LLaMA-3.1-70b exhibits the best performance. The effectiveness of BERT-base methods may be attributed to the strengths of encoder-based models like BERT in classification tasks over decoder-based models. Utilizing BERT-based models instead of LLMs enhances the speed and accessibility of HintRank, reducing computational demands. Figure 12 shows that the accuracy improves as the rank difference between the hints increases, indicating it is harder to correctly order hints with closer ranks.
6. Conclusions
@@
Automatic hint generation is a task closely related to question answering and question generation problems. Effective hints can constitute a useful mechanism for engaging users in the information seeking process. In this paper, we introduced the first manually created dataset for hint generation and ranking. We also presented a new lightweight method for ranking and evaluating hints. To demonstrate the effectiveness of our dataset, we conducted experiments where humans attempted to answer questions with and without the use of hints. The results confirm that the hints are of sufficient quality to assist users. We then finetuned LLMs using our dataset, prompting them to generate new hints for different questions. The high quality of the generated hints indicates that our dataset is well-suited for finetuning LLMs for the HG task. We also finetuned BERT and LLMs on WikiHint for the task of hint ranking and evaluated their performance. The results reveal that encoder-based models outperform decoder-based models in hint ranking.
@@
In the future, we plan to generate personalized hints tailored to the knowledge of askers. The main challenge here will be to develop appropriate datasets and solutions for user profiling.